Lightspeed Systems
All Systems Operational

About This Site

We know you rely on Lightspeed Systems solutions to do Amazing things so we continuously monitor our services internally as well as through 3rd party services. If there are any interruptions in services, updates will be posted here.

Classroom Orchestrator ? Operational
Launch Dashboard Operational
Mobile Device Manager ? Operational
Rocket ? Operational
Hosted Mobile Filter Operational
LSAccess.me Operational
Lightspeed Systems Community Site / Web Site Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Sep 24, 2016
Completed - The scheduled maintenance has been completed.
Sep 24, 00:00 PDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 23, 22:01 PDT
Scheduled - We will be performing database maintenance starting on Friday, September 23rd at 10 pm PST and continuing until Saturday, September 24 at 12 am PST.
Mobile Manger, Classroom Orchestrator, and the Launch dashboard will be briefly unavailable during the duration of this work.
Sep 21, 12:01 PDT
Sep 22, 2016

No incidents reported.

Sep 21, 2016
Resolved - Traffic levels have returned to normal and command requests should be processed normally.
Sep 21, 13:40 PDT
Monitoring - We are currently observing a very high load of device command requests, and this high load may cause delays in devices receiving commands.
Sep 21, 09:27 PDT
Sep 20, 2016

No incidents reported.

Sep 19, 2016
Resolved - Our ISP has closed this issue, and we are no longer seeing high latency or lost packets.
Sep 19, 16:05 PDT
Update - We have received word from our ISP that they estimate the time to repair is 2 hours.
Sep 19, 15:18 PDT
Monitoring - The NOC Team has observed a service disruption impacting several hosted services and is currently working to resolve the issue. We will update this page as more information becomes available. We appreciate your patience and apologize for the disruption.

It appears to be related to an outage from our Internet Service Provider, so we will monitor the situation closely and update when connectivity is regained.
Sep 19, 11:50 PDT
Sep 18, 2016
Resolved - Our upstream provider has closed their ticket on this incident, and all services appear to be operating normally.
Sep 18, 12:13 PDT
Monitoring - We are observing normal traffic flows at this time, but will continue to monitor until our upstream provider closes their ticket on this issue.
Sep 18, 10:03 PDT
Identified - Our upstream provider is reporting fiber outages that are impacting the global availability of these two services. Some customers may not be able to reach these services. Our provider currently reports that they expect the issues to be resolved by 11 AM PDT.
Sep 18, 08:30 PDT
Sep 17, 2016

No incidents reported.

Sep 16, 2016

No incidents reported.

Sep 15, 2016

No incidents reported.

Sep 14, 2016

No incidents reported.

Sep 13, 2016

No incidents reported.

Sep 12, 2016

No incidents reported.

Sep 11, 2016

No incidents reported.

Sep 10, 2016

No incidents reported.